# Hasan Fallahgoul

### v2

@@ 3 Kernel Approximation Breakdown
Having established the theory-practice disconnect in Section 2, I now prove rigorously that standardization fundamentally alters the kernel approximation properties that justify RFF methods. This breakdown explains why high-dimensional methods cannot achieve their claimed theoretical properties and must rely on simpler mechanisms.
@@
where $(\omega,b)\sim\mathcal{N}(0,\gamma^{2}I_{K})\times\text{Uniform}[0,2\pi]$.
Then:
@@ 4.1 Minimax Risk Framework
I conduct the minimax analysis in the finite–dimensional random-feature space by representing each candidate predictor as $f_{\omega}(x)=\omega^{\top}z(x)$. This inner-product form serves three key purposes. (i) It mirrors the reproducing-kernel expansion of Gaussian-kernel regression: $z(x)$ collects the Random Fourier Features and $\omega$ specifies their linear weights, keeping the setup directly comparable to the kernel methods analysed by Kelly et al. (2024). (ii) Expressing the function class through the constrained parameter vector $\omega\in\mathbb{B}_{2}^{P}(B)$ converts an infinite-dimensional functional problem into a finite linear one, enabling PAC- and information-theoretic risk bounds via standard packing, Kullback-Leibler (KL), and Fano arguments. (iii) The factorisation $\omega^{\top}z(x)$ neatly separates the learner-controlled parameters ($\omega$) from the data-driven randomness ($z(x)$), a separation that is crucial for deriving worst-case (minimax) prediction-error lower bounds while allowing probabilistic assumptions on the covariate distribution.
@@
The in-expectation bounds (parts (a) of Theorems 4.1 and 4.2) apply directly to the high-dimensional regime $P\gg T$ and provide fundamental limits on learning performance averaged over all possible feature realizations and datasets. The high-probability bounds (parts (b)) require $T\geq C_{0}P$ for technical reasons related to matrix concentration, making them inapplicable when $P\gg T$. However, the in-expectation bounds suffice to establish information-theoretic impossibility in practical high-dimensional scenarios.
@@ A.2.1 Step 1: Integrability and Almost-Sure Convergence
$\hat{\sigma}^{2}:=\frac{2}{T}\sum_{t=1}^{T}\cos^{2}(\omega^{\top}x_{t}+b)=1+S_{T},\quad S_{T}:=\frac{1}{T}\sum_{t=1}^{T}\cos(2\omega^{\top}x_{t}+2b)$
Since $|h|\leq 2\hat{\sigma}^{-2}$, integrability of $h$ follows once we show $\mathbb{E}[\hat{\sigma}^{-2}]<\infty$. Lemma A.1 proves this claim.
Using $\mathbb{P}(\hat{\sigma}^{-2}>u)=\mathbb{P}(\hat{\sigma}^{2}<u^{-1})$, we obtain:
@@
Since the variables $h(\omega_{i},b_{i})$ are i.i.d. with finite mean, Kolmogorov’s strong law yields:
@@ A.2.2 Step 2: Training-Set Dependence
The proof proceeds by perturbation analysis. We introduce a scaling factor $\alpha\geq 1$ for the training set and demonstrate that the derivative of the expectation with respect to $\alpha$ is non-zero at $\alpha=1$.
Let the training set be scaled by a parameter $\alpha$, denoted $\mathcal{T}(\alpha)=\{\alpha x_{1},...,\alpha x_{T}\}$. The function inside the expectation becomes a function of $\alpha$:
$h(\omega,b;\alpha)=\frac{2\cos(\omega^{\top}x+b)\cos(\omega^{\top}x^{\prime}+b)}{1+\frac{1}{T}\sum_{t=1}^{T}\cos(2\alpha\omega^{\top}x_{t}+2b)}=\frac{N(\omega,b)}{D(\omega,b;\alpha)}$
Our objective is to prove that for a generic, non-trivial training set, $\frac{\partial}{\partial\alpha}\mathbb{E}[h(\omega,b;\alpha)]\big{|}_{\alpha=1}\neq 0$.
The interchange of differentiation and expectation, $\frac{\partial}{\partial\alpha}\mathbb{E}[h]=\mathbb{E}[\frac{\partial h}{\partial\alpha}]$, is permitted by the Dominated Convergence Theorem if $|\frac{\partial}{\partial\alpha}h(\omega,b;\alpha)|$ is bounded by an integrable function $g(\omega,b)$ for $\alpha$ in a neighborhood of 1. The derivative’s magnitude is bounded by:
$\left|\frac{\partial h(\omega,b;\alpha)}{\partial\alpha}\right|\leq\frac{4}{T\cdot[D(\omega,b;\alpha)]^{2}}\sum_{t=1}^{T}|\omega^{\top}x_{t}|$
A formal justification requires showing that the expectation of this bound is finite. This relies on extending the logic of the small-ball estimates used in the proof of Theorem 1(a) to demonstrate that $\mathbb{E}\left[[D(\omega,b;\alpha)]^{-2}\sum_{t}|\omega^{\top}x_{t}|\right]<\infty$. Assuming this standard but technical verification holds, the interchange is justified.
Using the chain rule, the partial derivative of the integrand with respect to $\alpha$ is:
$\frac{\partial h(\omega,b;\alpha)}{\partial\alpha}$ $=-N(\omega,b)\cdot[D(\omega,b;\alpha)]^{-2}\cdot\frac{\partial}{\partial\alpha}\left(\frac{1}{T}\sum_{t=1}^{T}\cos(2\alpha\omega^{\top}x_{t}+2b)\right)$
$=-N(\omega,b)\cdot[D(\omega,b;\alpha)]^{-2}\cdot\left(-\frac{1}{T}\sum_{t=1}^{T}\sin(2\alpha\omega^{\top}x_{t}+2b)\cdot(2\omega^{\top}x_{t})\right)$
$=\frac{2N(\omega,b)}{T\cdot[D(\omega,b;\alpha)]^{2}}\sum_{t=1}^{T}(\omega^{\top}x_{t})\sin(2\alpha\omega^{\top}x_{t}+2b)$
We take the expectation of the derivative and evaluate it at $\alpha=1$:
$\frac{\partial}{\partial\alpha}\mathbb{E}[h]\bigg{|}_{\alpha=1}=\mathbb{E}\left[\frac{2N(\omega,b)}{T\cdot[D(\omega,b;1)]^{2}}\sum_{t=1}^{T}(\omega^{\top}x_{t})\sin(2\omega^{\top}x_{t}+2b)\right]$
The expectation of a function over a symmetric domain is zero if the function is odd with respect to the variable of integration. Here, the integration is over $\omega\sim\mathcal{N}(0,\gamma^{2}I_{K})$ and $b\sim\text{Uniform}[0,2\pi]$. The distribution of $\omega$ is symmetric with respect to the origin. However, the integrand does not possess the required odd symmetry under the transformation $\omega\to-\omega$ that would guarantee the expectation vanishes. The presence of the data-specific vectors $\{x_{t},x,x^{\prime}\}$ and the phase variable $b$ breaks any simple symmetries. For a generic (i.e., not pathologically constructed) choice of data vectors, the complex interplay of terms will not create the perfect cancellation required for the integral to be exactly zero. Thus, the derivative is non-zero.
Since the derivative of the limiting kernel with respect to the training set scaling is non-zero, the limiting kernel $k_{std}^{*}$ must depend on the training set $\mathcal{T}$.
A.3 Supporting Lemmas

### v3

@@ 4 Kernel Approximation Breakdown
Having established the theory-practice disconnect in Section 3, I now prove rigorously that standardization fundamentally alters the kernel approximation properties that justify RFF methods.
@@
where $(\omega,b)\sim\mathcal{N}(0,\gamma^{2}I_{K})\times\text{Uniform}[0,2\pi]$. Then:
@@
Corollary 4.1 (Kernel Breakdown for the Two-Feature RFF Variant).
Let the conditions of Theorem 4.1 hold. The standardized kernel derived from the feature pairs $[\sin(\omega^{\top}x),\cos(\omega^{\top}x)]$, as used in Kelly et al. (2024), also converges to a data-dependent limit $k_{\text{std}}^{*}(\mathcal{T})$ that is not, in general, the target Gaussian kernel $k_{G}$.
@@ 5.1 Minimax Risk Framework
I conduct the minimax analysis in the finite–dimensional random-feature space by representing each candidate predictor as $f_{\omega}(x)=\omega^{\top}z(x)$. This inner-product form serves three key purposes. First, tt mirrors the reproducing-kernel expansion of Gaussian-kernel regression: $z(x)$ collects the Random Fourier Features and $\omega$ specifies their linear weights, keeping the setup directly comparable to the kernel methods analysed by Kelly et al. (2024). Second, expressing the function class through the constrained parameter vector $\omega\in\mathbb{B}_{2}^{P}(B)$ converts an infinite-dimensional functional problem into a finite linear one, enabling PAC- and information-theoretic risk bounds via standard packing, Kullback-Leibler (KL), and Fano arguments. Third, the factorisation $\omega^{\top}z(x)$ neatly separates the learner-controlled parameters ($\omega$) from the data-driven randomness ($z(x)$), a separation that is crucial for deriving worst-case (minimax) prediction-error lower bounds while allowing probabilistic assumptions on the covariate distribution.
@@
The in-expectation bounds (parts (a) of Theorems 5.1 and 5.2) apply directly to the high-dimensional regime $P\gg T$ and provide fundamental limits on learning performance averaged over all possible feature realizations and datasets. The high-probability bounds (parts (b)) require $T\geq C_{0}P$ for technical reasons related to matrix concentration, making them inapplicable when $P\gg T$. However, the in-expectation bounds suffice to establish information-theoretic impossibility in practical high-dimensional scenarios.
@@ A.2.1 Step 1: Integrability and Almost-Sure Convergence
$\hat{\sigma}^{2}(\omega,b):=\frac{2}{T}\sum_{t=1}^{T}\cos^{2}(\omega^{\top}x_{t}+b)=1+S_{T},\quad S_{T}:=\frac{1}{T}\sum_{t=1}^{T}\cos(2\omega^{\top}x_{t}+2b).$ (A.1)
From the construction in Assumption 3.2, we have $|\cos(\cdot)|\leq 1$, which gives:
$|h(\omega,b)|\leq\frac{2}{1+\frac{1}{T}\sum_{t=1}^{T}\cos(2\omega^{\top}x_{t}+2b)}=:\frac{2}{\hat{\sigma}^{2}(\omega,b)}$ (A.2)
where $\hat{\sigma}^{2}(\omega,b)$ represents the empirical variance term from the standardization procedure.
The critical step is showing $\mathbb{E}[\hat{\sigma}^{-2}(\omega,b)]<\infty$. The denominator $\hat{\sigma}^{2}(\omega,b)$ can become small only when:
$\frac{1}{T}\sum_{t=1}^{T}\cos(2\omega^{\top}x_{t}+2b)\approx-1.$ (A.3)
This requires the system of equations:
$2\omega^{\top}x_{t}+2b\equiv\pi\pmod{2\pi},\quad t=1,\ldots,T.$ (A.4)
Under Assumption 3.4 (Affine Independence), the matrix $A=[x_{1}^{\top}\cdots x_{T}^{\top};1\cdots 1]^{\top}\in\mathbb{R}^{T\times(K+1)}$ has full column rank $T$. This ensures that system (A.4) defines a measure-zero set in the $(\omega,b)$ parameter space.
By Lemma A.1 with the affine independence condition from Assumption 3.4, there exists $C_{T}<\infty$ such that for every $\epsilon\in(0,1)$:
$\mathbb{P}(\hat{\sigma}^{2}\leq\epsilon)\leq C_{T}\epsilon^{T/2}.$ (A.5)
This yields:
@@
[…]
for $T>2$, as guaranteed by typical sample sizes in financial applications under Assumption 3.1.
Since $h(\omega_{i},b_{i})$ are i.i.d. with $\mathbb{E}[|h|]\leq 2\mathbb{E}[\hat{\sigma}^{-2}]<\infty$, and the regularity conditions in Assumption 3.3 ensure proper measurability, the Strong Law of Large Numbers applies:
@@ A.2.2 Step 2: Training-Set Dependence
I prove $k^{*}_{std}\neq k_{G}$ by demonstrating that these kernels possess incompatible mathematical structures. Specifically, $k_{G}$ is shift-invariant while $k^{*}_{std}$ is not.
The theoretical Gaussian kernel has the shift-invariant structure:
$k_{G}(x,x^{\prime})=\exp\left(-\frac{\gamma^{2}}{2}\|x-x^{\prime}\|^{2}\right)$ (A.10)
This kernel depends only on the relative distance $\|x-x^{\prime}\|$ between query points, ensuring that for any translation vector $c\in\mathbb{R}^{K}$:
$k_{G}(x+c,x^{\prime}+c)=k_{G}(x,x^{\prime}).$ (A.11)
This shift-invariance is a defining characteristic that underlies all theoretical guarantees of RFF methods.
The standardized limit kernel has the structure:
$k^{*}_{std}(x,x^{\prime}|\mathcal{T})=\mathbb{E}_{(\omega,b)}\left[\frac{N(x,x^{\prime},\omega,b)}{D(\mathcal{T},\omega,b)}\right]$ (A.12)
where:
$N(x,x^{\prime},\omega,b)$ $=2\cos(\omega^{T}x+b)\cos(\omega^{T}x^{\prime}+b)$ (A.13)
$D(\mathcal{T},\omega,b)$ $=1+\frac{1}{T}\sum_{t=1}^{T}\cos(2\omega^{T}x_{t}+2b).$ (A.14)
The critical observation is that the denominator $D(\mathcal{T},\omega,b)$ depends explicitly on the absolute positions of the training points $\{x_{t}\}_{t=1}^{T}$, not on their relative distances. This dependence structure is fundamentally incompatible with shift-invariance because:
  • Absolute position dependence: The term $\cos(2\omega^{T}x_{t}+2b)$ in the denominator depends on the coordinate values $x_{t}$ in the chosen coordinate system, not on pairwise distances $\|x_{i}-x_{j}\|$.
  • Fixed training set: When query points are shifted $x\to x+c$, the training set $\mathcal{T}=\{x_{1},\ldots,x_{T}\}$ remains fixed in the original coordinate system, so $D(\mathcal{T},\omega,b)$ is invariant under query point transformations.
  • Asymmetric transformation: Under the shift $x\to x+c,x^{\prime}\to x^{\prime}+c$, we have:
    $N(x+c,x^{\prime}+c,\omega,b)$ $=2\cos(\omega^{T}x+\omega^{T}c+b)\cos(\omega^{T}x^{\prime}+\omega^{T}c+b)$ (A.15)
[…]
This creates an asymmetric transformation where the numerator depends on the shift through $\omega^{T}c$ while the denominator remains fixed. For shift-invariance to hold, both numerator and denominator would need to transform consistently.
To prove the kernels are unequal, observe that:
$k^{*}_{std}(x+c,x^{\prime}+c|\mathcal{T})=\mathbb{E}_{(\omega,b)}\left[\frac{2\cos(\omega^{T}x+\omega^{T}c+b)\cos(\omega^{T}x^{\prime}+\omega^{T}c+b)}{1+\frac{1}{T}\sum_{t=1}^{T}\cos(2\omega^{T}x_{t}+2b)}\right].$ (A.17)
Under Assumption 3.2, $\omega^{T}c\sim\mathcal{N}(0,\gamma^{2}\|c\|^{2})$ for any $c\neq 0$. The transformation $\cos(\omega^{T}x+b)\mapsto\cos(\omega^{T}x+\omega^{T}c+b)$ introduces a random phase shift that changes the joint distribution of the numerator terms, while the denominator distribution remains unchanged.
Since there is no measure-preserving transformation that simultaneously maps the shifted numerator to the original numerator while keeping the denominator fixed, the expectations cannot be equal for generic choices of $x,x^{\prime},c,\mathcal{T}$.
Therefore, $k^{*}_{std}$ violates shift-invariance: $k^{*}_{std}(x+c,x^{\prime}+c|\mathcal{T})\neq k^{*}_{std}(x,x^{\prime}|\mathcal{T})$ for $c\neq 0$. Since $k_{G}$ is shift-invariant by construction, we conclude $k^{*}_{std}\neq k_{G}$.
A.3 Proof of Corollary 4.1
Proof of Corollary 4.1.
We prove that the two-feature RFF variant employed by Kelly et al. (2024) suffers the same kernel approximation breakdown under standardization as the standard RFF construction.
For each random draw $\omega_{i}\sim\mathcal{N}(0,\gamma^{2}I_{K})$, the two-feature variant generates the pair:
$s_{i}(x)=\sin(\omega_{i}^{\top}x),\quad c_{i}(x)=\cos(\omega_{i}^{\top}x).$ (A.18)
The key theoretical property is that these features satisfy the trigonometric identity:
$s_{i}(x)s_{i}(x^{\prime})+c_{i}(x)c_{i}(x^{\prime})$ $=\sin(\omega_{i}^{\top}x)\sin(\omega_{i}^{\top}x^{\prime})+\cos(\omega_{i}^{\top}x)\cos(\omega_{i}^{\top}x^{\prime})$ (A.19)
$=\cos(\omega_{i}^{\top}(x-x^{\prime})).$ (A.20)
Equation (A.19) deterministically yields the desired shift-invariant kernel approximation without the Monte Carlo randomness of the phase shift $b_{i}$ in standard RFF.
In practice, features are standardized using training sample statistics. Define:
$\hat{\sigma}^{2}_{s,i}$ $=\frac{1}{T}\sum_{t=1}^{T}s_{i}(x_{t})^{2}=\frac{1}{T}\sum_{t=1}^{T}\sin^{2}(\omega_{i}^{\top}x_{t})$ (A.21)
$\hat{\sigma}^{2}_{c,i}$ $=\frac{1}{T}\sum_{t=1}^{T}c_{i}(x_{t})^{2}=\frac{1}{T}\sum_{t=1}^{T}\cos^{2}(\omega_{i}^{\top}x_{t}).$ (A.22)
where the standardized features in (A.21) are:
$\tilde{s}_{i}(x)=\frac{s_{i}(x)}{\hat{\sigma}_{s,i}},\quad\tilde{c}_{i}(x)=\frac{c_{i}(x)}{\hat{\sigma}_{c,i}}.$ (A.23)
Thus, the standardized kernel estimator becomes:
$k^{(P)}_{std,2F}(x,x^{\prime})=\frac{1}{P}\sum_{i=1}^{P}\left[\tilde{s}_{i}(x)\tilde{s}_{i}(x^{\prime})+\tilde{c}_{i}(x)\tilde{c}_{i}(x^{\prime})\right].$ (A.24)
Using the half-angle formulas:
$\sin^{2}(\theta)$ $=\frac{1-\cos(2\theta)}{2}$ (A.25)
$\cos^{2}(\theta)$ $=\frac{1+\cos(2\theta)}{2}.$ (A.26)
We can rewrite the standardization factors as the following:
$\hat{\sigma}^{2}_{s,i}$ $=\frac{1}{T}\sum_{t=1}^{T}\sin^{2}(\omega_{i}^{\top}x_{t})=\frac{1}{2}-\frac{1}{2T}\sum_{t=1}^{T}\cos(2\omega_{i}^{\top}x_{t})$ (A.27)
$\hat{\sigma}^{2}_{c,i}$ $=\frac{1}{T}\sum_{t=1}^{T}\cos^{2}(\omega_{i}^{\top}x_{t})=\frac{1}{2}+\frac{1}{2T}\sum_{t=1}^{T}\cos(2\omega_{i}^{\top}x_{t}).$ (A.28)
By define $S_{T}(\omega)=\frac{1}{T}\sum_{t=1}^{T}\cos(2\omega^{\top}x_{t})$, we have
$\hat{\sigma}^{2}_{s}(\omega)=\frac{1-S_{T}(\omega)}{2},\quad\hat{\sigma}^{2}_{c}(\omega)=\frac{1+S_{T}(\omega)}{2}.$ (A.29)
Define the two-feature standardized kernel function as the following:
$h_{2F}(\omega)=\frac{\sin(\omega^{\top}x)\sin(\omega^{\top}x^{\prime})}{\hat{\sigma}^{2}_{s}(\omega)}+\frac{\cos(\omega^{\top}x)\cos(\omega^{\top}x^{\prime})}{\hat{\sigma}^{2}_{c}(\omega)}.$ (A.30)
We need to establish $\mathbb{E}[|h_{2F}(\omega)|]<\infty$. Since $|\sin(\cdot)|\leq 1$ and $|\cos(\cdot)|\leq 1$:
$|h_{2F}(\omega)|\leq\frac{1}{\hat{\sigma}^{2}_{s}(\omega)}+\frac{1}{\hat{\sigma}^{2}_{c}(\omega)}.$ (A.31)
Note that $\hat{\sigma}^{2}_{s}(\omega)$ approaches zero when $S_{T}(\omega)\to 1$, which requires:
$\cos(2\omega^{\top}x_{t})\approx 1\text{ for all }t=1,\ldots,T.$ (A.32)
This is equivalent to the system:
$2\omega^{\top}x_{t}\equiv 0\pmod{2\pi},\quad t=1,\ldots,T.$ (A.33)
Under Assumption 3.4 (affine independence), the augmented matrix $[x_{1}\cdots x_{T}]^{\top}$ has full rank. By a modification of Lemma A.1 for the doubled angles, there exists $C^{\prime}_{T}<\infty$ such that for $\varepsilon\in(0,1/2)$:
$\mathbb{P}\left(\hat{\sigma}^{2}_{s}(\omega)\leq\varepsilon\right)$ $\leq C^{\prime}_{T}(2\varepsilon)^{T/2}$ (A.34)
$\mathbb{P}\left(\hat{\sigma}^{2}_{c}(\omega)\leq\varepsilon\right)$ $\leq C^{\prime}_{T}(2\varepsilon)^{T/2}.$ (A.35)
Following the argument in equation (A.6)of Theorem 4.1:
[…]
$=1+\frac{2^{T/2+1}C^{\prime}_{T}}{T-2}<\infty$ (A.39)
for $T>2$. The same bound holds for $\mathbb{E}[1/\hat{\sigma}^{2}_{c}]$.
Since $h_{2F}(\omega_{i})$ are i.i.d. random variables with $\mathbb{E}[|h_{2F}|]<\infty$, the Strong Law of Large Numbers yields:
$k^{(P)}_{std,2F}(x,x^{\prime})\xrightarrow[P\to\infty]{a.s.}k^{*}_{std,2F}(x,x^{\prime}|\mathcal{T}):=\mathbb{E}_{\omega}[h_{2F}(\omega)].$ (A.40)
The limiting kernel has the explicit form:
$k^{*}_{std,2F}(x,x^{\prime}|\mathcal{T})=\mathbb{E}_{\omega}\left[\frac{2\sin(\omega^{\top}x)\sin(\omega^{\top}x^{\prime})}{1-S_{T}(\omega)}+\frac{2\cos(\omega^{\top}x)\cos(\omega^{\top}x^{\prime})}{1+S_{T}(\omega)}\right].$ (A.41)
Note that this kernel depends explicitly on the training set $\mathcal{T}=\{x_{1},\ldots,x_{T}\}$ through $S_{T}(\omega)$. Crucially, since $\hat{\sigma}^{2}_{s}(\omega)\neq\hat{\sigma}^{2}_{c}(\omega)$ in general (they are equal only when $S_{T}(\omega)=0$), we have:
$\frac{\sin(\omega^{\top}x)\sin(\omega^{\top}x^{\prime})}{\hat{\sigma}^{2}_{s}(\omega)}+\frac{\cos(\omega^{\top}x)\cos(\omega^{\top}x^{\prime})}{\hat{\sigma}^{2}_{c}(\omega)}\neq\cos(\omega^{\top}(x-x^{\prime})).$ (A.42)
Under the shift transformation $x\mapsto x+c$, $x^{\prime}\mapsto x^{\prime}+c$ for $c\in\mathbb{R}^{K}$:
$h_{2F}(\omega;x+c,x^{\prime}+c)$ $=\frac{\sin(\omega^{\top}x+\omega^{\top}c)\sin(\omega^{\top}x^{\prime}+\omega^{\top}c)}{\hat{\sigma}^{2}_{s}(\omega)}$ (A.43)
$\quad+\frac{\cos(\omega^{\top}x+\omega^{\top}c)\cos(\omega^{\top}x^{\prime}+\omega^{\top}c)}{\hat{\sigma}^{2}_{c}(\omega)}.$ (A.44)
The denominators $\hat{\sigma}^{2}_{s}(\omega)$ and $\hat{\sigma}^{2}_{c}(\omega)$ remain unchanged as they depend only on the fixed training set $\mathcal{T}$. However, the numerators transform non-trivially through the phase shift $\omega^{\top}c$.
For shift-invariance to hold, we would need:
$k^{*}_{std,2F}(x+c,x^{\prime}+c|\mathcal{T})=k^{*}_{std,2F}(x,x^{\prime}|\mathcal{T}).$ (A.45)
This would require the expectation to be invariant under the transformation induced by $\omega^{\top}c\sim\mathcal{N}(0,\gamma^{2}\|c\|^{2})$. Since the numerator and denominator transform asymmetrically (numerator changes, denominator fixed), this invariance cannot hold for generic $x,x^{\prime},c,\mathcal{T}$.
Since the Gaussian kernel $k_{G}(x,x^{\prime})=\exp(-\gamma^{2}\|x-x^{\prime}\|^{2}/2)$ is shift-invariant while $k^{*}_{std,2F}$ is not, we conclude:
$k^{*}_{std,2F}(x,x^{\prime}|\mathcal{T})\neq k_{G}(x,x^{\prime}).$ (A.46)
Therefore, standardization destroys the kernel approximation properties of the two-feature RFF variant, despite its theoretical advantage of deterministically achieving $\cos(\omega^{\top}(x-x^{\prime}))$ in the unstandardized case. ∎
A.4 Supporting Lemmas